Show the output produced by each of the following program fragments. Assume that I, j, k...

  1. Show the output produced by each of the following program fragments. Assume that I, j, k are int variables
    1. i = 7; j = 2;

printf(“%d %d ”, i / j, i % j);

  1. i = 4; j = 3;

printf(“%d”, (i + 10 ) % j);

  1. i = 7; j = 8; k = 9;

printf("%d”, (i + 10) % k / j);

  1. Show the output produced by each of the following program fragments. Assume that I, j, k are int variables
    1. i = 7; j = 8;

i *= j + 3;

printf(“%d %d”, i, j);

  1. i = j = k = 1;

i *= j *= k;

printf(“%d %d %d”, i, j, k);

  1. Show the output produced by each of the following program fragments. Assume the I, j, k are int variables.
    1. i = 4;

j = ++i * 3 – 2;

printf(“%d %d” i, j)

  1. i = 3;

j = 3 – 2 * i++;

printf(“%d %d”, i, j);

  1. i = 2; j = 5;

printf(“%d ”, i++ - ++j);

printf(“%d %d”, i, j);
Answer #1

For the first part:

a. 3 1

(7/2=3(gives the rounded value to the lowest whole number, so 3.5 becomes 3)).

(7%2 dividing 7 with 2 will give you a remainder of 1. So % checks the remainder also known as modulo).

b. 2

(14 divided by 3 will give us a remainder of 2) Therefore 2. to check 3*4=12, So 14-12=2.(Just another way).

c. 1

(17 divided by 9 will give a remainder of 8. So 8 divided by 8 will give a quotient of 1. Therefore answer is 1.)

For the second part:

a. 77 8

This expression is also equal to i=i*(j+3). Therefore 7*(8+3)=7*11=77; and j gets printed the same so 8.

b. 1 1 1

This expression will start from the end. so k remains the same=1 j=j*k=1 and i=i*j=1;

Suppose, instead of 1 we took 4, answer would be k=4 j=j*k, so j=4*4=16 and i=i*j, so i=4*16=64.

For the Third part:

a. 5 13

Here, i is a pre-incrementation operation. So, first it will increase the value, then it will operate. So i becomes 4+1=5, and j=(5*3)-2=13.

b. 4 -3

Here, i is a post-incrementation operation. So, it will increase its value after the operation is over. So, the expression for j=3-2*3=-3 and then i will increase its value, therefore i=4;

c. -4

i is post incrementation and j is pre-incrementation. So, j will increase its value while the arithmetic operation is taking place, whereas i will increase after the arithmentic operation is over. So, 2-6=-4

3 6

Now, since the operation is over i will increase its value by 1, j has already done it during the expression so it remains the same

i=(2+1) and j=6.